OpenStudy (anonymous):

[6.05] Your piggy bank has a total of 43 coins in it; some are dimes and some are nickels. If you have a total of $3.60, how many nickels and how many dimes do you have? Show all of your work .

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

ok give me a variable for dimes and one for nickels

Mehek (mehek14):

@kaylamarie081 ?

OpenStudy (anonymous):

okay. um d= dimes and n= nickels

Mehek (mehek14):

alright so d + n = 43 coins in total

Mehek (mehek14):

d = 10 cents n = 5 cents 10d + 5n = 360

Mehek (mehek14):

so 10d + 5n = 360 d + n = 43

Mehek (mehek14):

10d + 5n = 360 10d + 10n = 430 <---- multiply by 10

Mehek (mehek14):

subtract both -5n = -70

Mehek (mehek14):

n = 14 14 nickels

Mehek (mehek14):

substitute in the equation d + 14 = 43 d = 29

Mehek (mehek14):

so 14 nickels and 29 dimes 14 * 0.05 = 0.7 29 * 0.1 = 2.9 0.7 + 0.9 = 3.6

Mehek (mehek14):

so the answer is 14 nickels and 29 dimes @kaylamarie081

Mehek (mehek14):

did you get it?

Mehek (mehek14):

@kaylamarie081 ????

OpenStudy (anonymous):

yes thank you. sorry it took awhile i was trying to figure it out on my own for a mintue.